How do I calculate the p value , when doing calculating the chi square value in genetics , please include formula
Answer #1

Chi square = (observed frequency - expected frequency)^2 / expected frequency = x^2

Summation chi sqaure = add all x^2 values = calculated chi square value

Degree of freedom = number of groups - 1

Now, you have to find out critical chi sqaure value using the chi square table.

You have to check critical value at a p value of 5% or 0.05.

Suppose, calculated chi sqaure value is 0.234 and degress of freedom is 3. Use , a chi sqaure table and find out where number 0.234 will lie when degree of freedom is 3. 0.234 will lie between 1.212 and 2.366 which a p value of 0.75 and 0.50 respectively.

To accept or reject chi square, find chi sqaure value at p 0.05 and degree of freedom 3. It is 7.81 = critical value.

Our critical value is more than calculated value. So, we will accept our null hypothesis.

Please rate.